I really need help solving this problem using Cramer's rule {x-y=2
y-x+4=0
I don't understand how to do it when there's an extra number involved (like the 4 in this case)

Answer by jsmallt9(3759)

Just use algebra to get the second equation into the form you need for Cramer's rule. Subtracting 4 from each side we get:
y - x = -4
You also want the x in front. So change it to addition (so you can change the order of the terms):
y + (-x) = -4
Now put the x in front:
(-x) + y = -4

Now your system is:
x - y = 2
-x + y = -4

You should be able to use Cramer's rule now.